# Motion Capture from Inertial and Vision Sensors

### v1

@@ 4.1 Qualitative Results
We show the qualitative results of our annotated 3D SMPL mesh and the corresponding re-projected full-body 2D joints from different viewpoints in Figure 6 and Figure 7. Benefiting from the multi-view vision sensors and full-body inertial devices, our pipeline can recover the whole-body human motion and global position under various environmental conditions.
4.2 Multi-modal Human Motion Capture
In this subsection, we perform a baseline analysis of multi-modal human motion capture on our MINIONS dataset and explore the potential for consumer-affordable motion capture using a monocular camera and a minimal number of IMUs.
Dataset. We perform all experiments of this subsection on our MINIONS-S dataset. We divide our dataset into a training set, a validation set, and a testing set by actors. The training set contains 12 actors with 2.2 million frames. The validation set has three actors with 0.9 million frames. The testing set has five actors with 1.4 million frames. To evaluate the performance of motion capture, we use the following metrics: the mean global rotation error $\mu_{glo}$ of all body joints (in degree), the MPJPE of all body joints (in mm), and the average jerk of body joints [57] (in $10^{2}m/s^{3}$).
Baseline Model. To explore consumer-affordable multi-modal motion capture, we utilize the multi-model motion capture model SparseNet to explore the supplementary features from inertial and vision sensors in this subsection. SparseNet is designed based on the real-time monocular motion capture model ROMP [46] and the sparse IMUs-based motion capture model Transpose [57]. This incremental baseline model enables us to emphasize the distinct contributions of each modality, without being obscured due to differences in model architecture and training settings. Please refer to the supplementary materials for the details of our baseline model.
Implementation Details. During training, the input frames of videos are resized to $512\times 512$ and the batch size is set as 512. The model is well-trained for 200 epochs using Adam optimizers with a learning rate of 1e-3. All training and test processes run on an NVIDIA GTX 3090 GPU.
@@ 4.2 Multi-modal Human Motion Capture
Table 2: Comparation between IMUs-based, monocular vision-based, and multi-modal human motion capture. #IMUs: the number and placement of IMUs used in the algorithms. In detail, 4 means {LeftHand, RightHand, Head, Hip}, 6 means {LeftForeArm, RightForeArm, LeftLeg, RightLeg, Head, Hip}, 6 means {LeftHand, RightHand, LeftFoot, RightFoot, Head, Hip}, 8 means {LeftHand, RightHand, LeftForeArm, RightForeArm, LeftFoot, RightFoot, Head, Hip}, 10 means {LeftHand, RightHand, LeftForeArm, RightForeArm, LeftFoot, RightFoot, LeftLeg, RightLeg, Head, Hip}. #Cams: the number of camera views used in the algorithms. $\mu_{glo}$ and $\sigma_{glo}$: the mean and variance global rotation error of all body joints in degrees. MPJPE: the mean Euclidean distance between the predicted 3D joint positions and the corresponding ground truth joint positions. Jitter: the average jerk of body joints.
#IMUs #Cams $\mu_{glo}$$\downarrow$ [deg] $\sigma_{glo}$$\downarrow$ [deg] MPJPE $\downarrow$ [mm] Jitter $\downarrow$ [$10^{2}m/s^{3}$]
IMUs-based 4 0 14.87 12.27 70.13 10.71
6 0 11.74 8.79 57.31 11.17
6 0 13.58 10.89 64.64 11.03
8 0 11.09 8.11 53.25 11.06
10 0 11.06 8.19 52.71 11.31
Vision-based 0 1 10.31 7.10 45.44 51.19
Multi-modal 4 1 9.82 6.93 41.93 2.82
6 1 9.23 6.22 39.87 2.24
6 1 9.41 6.61 41.91 2.82
8 1 8.89 6.05 39.78 2.69
10 1 8.82 5.98 39.24 2.59
Comparison. We train and evaluate the baseline model based on different settings on our MINIONS dataset, including IMUs-based, monocular vision-based, and multi-modal human motion capture. Following previous works, we apply $\mu_{glo}$ (mean global rotation error), $\sigma_{glo}$ (variance of global rotation error), and MPJPE (Euclidean distance error) to measure the mean error between prediction and ground truth in Euclidean space and angular space. Additionally, we use the Jitter to measure the average jerk of body joints. Our experimental results are detailed in Table 2. The results from our experiments show that motion capture using the monocular camera is frequently hampered by instability resulting from blurriness or occlusion and leads to large Jitter. In contrast, motion capture systems that rely on IMUs are prone to larger errors in global rotations and Euclidean distance, attributable to the inherent unreliability of the IMUs. However, an integrated system combining just four to six IMUs with a monocular camera can successfully ensure stable motion capture with minimal error and virtually no jitter. Furthermore, our experimental results show that a motion capture system equipped with a monocular camera and more than eight IMUs is unnecessary. They provide negligible improvements in global rotation error and Euclidean distance error, while leading to increased jitter and equipment cost.
@@
Figure 8: Visualization comparisons among the vision-based, IMUs-based, and multi-modal human motion capture. The vertexes are colored by the distances to the ground truth positions.
@@
Visualization. To facilitate a more intuitive comparison, we provide visualization results of vision-based, IMUs-based, and multi-modal motion capture in Figure 8. The vertexes of the 3D human body are colored by the distances between the prediction and the ground truth positions. Additionally, we present the global pose error and positional translation for a motion capture sequence in Figure 9.
Figure 9: Visualization. (a): Average angular error (in degree) over sequences. (b): Average translation error (in mm) over sequences.
The results reveal that the motion capture based on IMUs tends to exhibit a gradual increase in both global pose error and global positional translation over time. However, the integration of a monocular camera and the IMUs can significantly reduce global pose error and also mitigate the inherent global positional drift caused by IMUs. Please refer to the supplementary materials for more visualization.
4.3 Benchmarks on other Tasks
@@
In this subject, we evaluate mainstream baselines and benchmarks on various research tasks including 2D-to-3D pose estimation [36, 40, 63] and Fine-grained Video Action Recognition [55, 26, 52] , which facilitate more comprehensive explorations of our dataset.
Table 3: 2D-to-3D pose estimation. #Num: the receptive field of input frames.
Method #Num MPJPE $\downarrow$ (mm) PA-MPJPE $\downarrow$ (mm)
@@
2D-to-3D pose estimation. The 2D-to-3D pose estimation task aims to accurately predict the location of 3D joints from the detected 2D human joints, which is a specific type of 3D pose estimation task. The challenge is that depth information is lost in the projection from 3D joints to 2D joints, making the reverse process of deducing 3D joints from 2D observations ill-posed. In Table 3, we evaluate the accuracy of the mainstream method. MPJPE is applied for measuring the average Euclidean distance between the ground truth and predicted joint positions, while PA-MPJPE (Procrustes-Aligned Mean Per Joint Position Error) is a similar metric that first aligns the predicted pose to the ground truth using a rigid transformation before computing the error. To explore the upper bounds of current mainstream methods, we use the groundtruth of 2D joints as input. The experimental results indicate that there is still room for improvement in this task.
@@
Fine-grained Video Action Recognition. Our dataset provides fine-grained action labels and motion videos from various views, which can be validated for applications in fine-grained video action recognition. As shown in Table 4, we test the accuracy of the mainstream video action understanding methods on our dataset. Following these methods, Top-1 and Top-5 are used to evaluate model performance. Top-1 is the model’s accuracy in predicting the most likely action, while Top-5 is the accuracy at which the true action is within the model’s top-five predictions. Compared to the mainstream video understanding datasets, e.g. Kinetics400 [12], NTU RGB+D 60 [43], and NTU RGB+D 120 [29], our dataset is more challenging since existing methods easily achieve Top-1 accuracy of more than 85% on the Kinetics400 dataset.
5 Discussions
5.1 Ethical Issues
The principal ethical consideration addressed in this paper revolves around the issue of privacy. It is our utmost priority to ensure the protection and confidentiality of the data pertaining to the individuals incorporated in our dataset. We recognize the importance of privacy and are committed to implementing robust measures to safeguard it. In addition, our dataset is not intended for indiscriminate distribution. It is specifically designed and will be exclusively disseminated for the purpose of academic and scholarly research. This will be enforced through the implementation of a rigorous licensing agreement, thereby ensuring that the data is used responsibly and in alignment with our ethical commitment to privacy.
Table 4: Fine-grained Video Action Recognition on our MINIONS dataset.
Method Top-1 Accuracy $\uparrow$ Top-5 Accuracy $\uparrow$
Val Test Val Test
TSN [53] 44.94% 47.74% 77.00% 82.87%
I3D [12] 56.82% 59.64% 88.74% 90.41%
R(2+1)D [49] 55.48% 55.59% 85.78% 88.78%
irCSN [48] 58.51% 61.82% 92.77% 92.67%
SlowFast [15] 53.93% 58.64% 83.82% 88.52%
TSM [28] 59.37% 54.31% 88.64% 85.61%
TPN [55] 64.57% 63.97% 92.11% 91.78%
VideoMAE [52] 73.75% 75.46% 96.01% 96.34%
UniFormerV2 [26] 75.88% 77.65% 96.87% 97.29%
5.2 Limitations
Although the experimental results of multi-modal motion capture show that the fusion of vision and inertial devices significantly improves motion capture accuracy, there still is much room for further research and improvement. Furthermore, the current multi-modal motion capture is not equipped to handle the motion capture of multiple actors, which necessitates the consideration of tracking and matching actors with multi-modal inputs. This constitutes a valuable and promising area of study that merits further scholarly exploration.
6 Conclusion

### v3

@@ III-E Dataset Statistics
The statistic on the number of frames for different action sequences reflects that most action sequences are smaller than 300 frames (6 seconds), and the longest one can be up to 1500 frames (50 seconds), which shows the complexity of human actions in life. We show some fine-grained action examples of our dataset in Figure 4. Moreover, we show the qualitative results of our annotated 3D SMPL mesh and the corresponding re-projected full-body 2D joints from different viewpoints in Figure 5 and Figure 6. Benefiting from the multi-view vision sensors and full-body inertial devices, our pipeline can recover the whole-body human motion and global position under various environmental conditions.
IV Multi-modal Human Motion Capture
In this section, we present a SparseNet framework to explore the supplementary features from a monocular video and sparse IMUs for lightweight and easy-to-use human motion capture. Before detailing the network structure, we first introduce the theoretical assumptions in the following section.
Figure 7: Overview of the multi-model motion capture framework SparseNet. (a) The visual branch is designed to regress the body shape parameters $\beta$ and prior pose distribution $p(\theta)$, and the corresponding bone direction distribution $p(d_{v}|\theta)$. (b) The Sparse IMUs Branch is designed to infer the comprehensive full-body rotation distribution $p(D_{I}|\theta)$ from noisy, sparse inertial signals. (c) The post-processing branch is designed to fusion data from inertial and vision sensors, and smooth the final pose $\theta$. (d) The global translation $t$ is computed as a PNP-like problem.
IV-A Theoretical Assumptions
When both camera and IMUs are available, the estimation of human joint rotation $\theta$ can be formulated as a probabilistic inference problem on the special orthogonal group $SO(3)$ and unit sphere $S^{2}$. To robustly integrate these heterogeneous sensor data and account for their respective uncertainties, we adopt a Bayesian fusion strategy. Following the previous theoretical foundation [13, 30, 17], we model the joint rotation $\theta$ as a latent variable with a prior distribution, and fuse multi-modal observations (visual and inertial). Specifically, the prior distribution of joint rotation is modeled as a Matrix Fisher distribution over $SO(3)$, denoted as $p(\theta)$. The visual branch estimates 3D bone directions distribution $S^{2}$ from monocular images, which can be formulated as directional observations following a von Mises-Fisher (VMF) distribution [30]:
$p(d_{v}|\theta)=\frac{1}{c(\kappa_{v})}\exp\left(\kappa_{v}l_{v}^{T}{\theta}^{T}d_{v}\right),$ (5)
where $l_{v}$ is the reference bone direction in the T-pose, and $\kappa_{v}$ is the concentration parameter reflecting confidence, $d_{v}={\theta}l_{v}$ is the current bone directional. Similarly, the IMU branch provides rotational observations $p(D_{I}|\theta)$. Assuming conditional independence of observations, the posterior distribution of joint rotation given both modalities is
$p(\theta|d_{v},D_{I})\propto p(\theta)\cdot p(d_{v}|\theta)\cdot p(D_{I}|\theta).$ (6)
This Bayesian formulation allows for flexible multi-sensor fusion, supporting both training and inference stages. It enables robust handling of sensor uncertainty, as the posterior distribution naturally reflects the reliability of each modality.
IV-B Network Structure
The overall architecture of SparseNet is shown in Figure 7. To effectively leverage the complementary strengths of vision and inertial sensing, we design a two-branch structure. The visual and sparse IMUs branch is designed to learn latent features from visual images and sparse inertial sensor signals. Then, the posterior fusion module aligns the pose distributions from both branches, resulting in a more accurate and smoother pose rotation distribution. Finally, we consider the problem of computing translation $t$ from camera parameters $(K,R|T)$, the SMPL [38] pose rotation $\theta$, and shape $\beta$ as a Perspective-N-Point (PNP) problem and solve it through the PNP solver.
Visual Branch is designed to reconstruct the 3D body shape $\beta$, prior pose distribution $p(\theta)$ and the bone direction distribution $p(d_{v}|\theta)$ of subjects from monocular RGB frames, as depicted in Fig. 7 (a). This branch addresses the limitations of IMU-only systems by providing rich visual context and accurate bone locations, which are crucial for mitigating global drift and recovering human body shapes. Let $\mathbf{V}=\left\{V_{i}\right\}_{i=1}^{L}$ denote the input video sequence, where each $V_{i}\in\mathbb{R}^{3\times H\times W}$ represents the $i$-th RGB frame with spatial resolution $H\times W$, and $L$ is the total number of frames. The visual branch employs a vision encoder $f_{ve}$, based on the Vision Mamba [73] backbone, to extract visual features $z_{v}$ from the input frames. These features are subsequently processed by a lightweight shape decoder $f_{sd}$ and a pose decoder $f_{pd}$ to regress the SMPL [38] body shape parameters $\beta$ and prior pose distribution $p(\theta)$. Besides, a bone decoder $f_{bd}$ is used for estimating the bone directions $d_{v}$ and the corresponding bone direction distribution $p(d_{v}|\theta)$ is computed according to Eq. 5. This architectural design enables the visual branch to robustly recover detailed 3D body shape and pose information from single-view RGB inputs, facilitating accurate full-body reconstruction.
@@ IV-B Network Structure
TABLE II: Comparison between IMUs-based, monocular vision-based, and multi-modal human motion capture. #IMUs: the number and placement of IMUs used in the algorithms. In detail, 4 means {LeftHand, RightHand, Head, Hip}, 6 means {LeftForeArm, RightForeArm, LeftLeg, RightLeg, Head, Hip}, 8 means {LeftHand, RightHand, LeftForeArm, RightForeArm, LeftFoot, RightFoot, Head, Hip}, 10 means {LeftHand, RightHand, LeftForeArm, RightForeArm, LeftFoot, RightFoot, LeftLeg, RightLeg, Head, Hip}. #Cams: the number of camera views used in the algorithms. $\mu_{glo}$ and $\sigma_{glo}$: the mean and variance global rotation error of all body joints in degrees. MPJPE: the mean Euclidean distance between the predicted 3D joint positions and the corresponding ground truth joint positions. Jitter: the average jerk of body joints.
#IMUs #Cams $\mu_{glo}\downarrow$ [deg] $\sigma_{glo}\downarrow$ [deg] MPJPE $\downarrow$ [$m$] Jitter $\downarrow$ $10^{2}m/s^{3}$]
PIP [65] 4 0 16.79 13.59 79.17 1.74
6 0 15.73 12.53 77.23 1.95
8 0 13.46 10.20 64.88 1.35
10 0 13.46 10.29 64.35 1.23
PNP [67] 4 0 15.94 12.35 74.93 1.25
6 0 14.53 11.67 71.39 1.40
8 0 12.52 10.38 57.75 1.69
10 0 12.06 10.21 55.16 1.54
IMUs-based 4 0 14.79 12.04 68.94 1.71
6 0 11.67 8.65 57.93 1.17
8 0 11.20 8.06 53.55 1.06
10 0 11.09 8.31 53.20 1.31
TokenHMR [15] 0 1 12.89 6.16 49.36 10.06
PromptHMR [60] 0 1 11.45 8.87 47.67 11.67
Vision-based 0 1 10.27 7.20 45.61 13.02
Multi-modal 4 1 9.75 6.94 41.53 1.79
6 1 9.20 6.19 39.99 1.57
8 1 8.86 5.98 39.39 1.70
10 1 8.81 5.95 39.43 1.63
Sparse IMUs Branch is designed to infer comprehensive full-body pose information from noisy, sparse inertial signals, as illustrated in Fig. 7 (b). This branch is crucial for providing high-frequency, occlusion-resistant motion cues that complement the visual modality, especially during fast movements or partial occlusions where visual information may be unreliable. Formally, we denote the sequence of real sparse inertial input signals as $\mathbf{I}=\left\{I_{i}\right\}_{i=0}^{L}$, where each $I_{i}\in\mathbb{R}^{NC}$ represents the $i$-th sample containing the global orientation and acceleration of bones in the global coordinate system. Here, $N$ is the number of inertial sensors, $C$ is the dimensionality of the inertial data. Initially, the sparse signals $I_{i}$ are processed by the joint mamba [12] backbone $f_{jm}$, which predicts the bone positions $d_{0:i}$ according to $d_{0:i}=f_{jm}(I_{0:i})$. Compared to directly predicting rotation and global translation, introducing the joint mamba backbone helps improve the stability of the model. Subsequently, leveraging the estimated bone positions $d_{0:i}$, the IMU mamba encoder, translation decoder, and rotation decoder are employed to estimate the global translation $t_{I}$, joint rotations $D_{I}$, and the corresponding rotation distribution $p(D_{I}|\theta)$. Note that the global translation $t_{I}$ is used for IMUs-based experiments in section V.
@@
Figure 8: Visualization. (a): Average angular error (in degrees) over sequences. (b): Average translation error (in $m$) over sequences. (c): Visualization comparisons among the IMUs-based, vision-based, and multi-modal human motion capture. The vertices are colored by the distances to the GT.
Post-processing Branch, as shown in Fig. 7 (c), is designed to integrate the pose distribution $p(\theta)$, the visual bone direction distribution $p(d_{v}|\theta)$, and the IMUs rotation distribution $p(D_{I}|\theta)$ from both the visual and IMUs branches, to estimate the fused pose distribution $p(\theta|d_{v},D_{I})$ as described in Eq. 6. This fusion module is central to our framework, as it dynamically weights the contributions from each sensor modality based on their estimated uncertainties, thereby achieving more robust and accurate pose estimation than either modality alone. After that, the smooth mamba backbone is applied for smoothing the final pose $\theta$. With the fusion pose estimation results, we consider the problem of computing translation $t$ as a PNP-like problem in Fig. 7 (d). With the predicted human shape $\beta$, pose $\theta$, 2D joints $P_{2d}$, and given camera parameters $K$, we can easily obtain the global transition $t$ by
$\arg\min_{t}\left|{P}_{2d}-K\mathcal{J}M({\theta},{\beta},t)\right|,$ (7)
where $K\mathcal{J}$ is a mapping matrix from SMPL mesh $M({\theta},{\beta},t)$ to 2D keypoints, which is pre-trained in [38].
V Experiments
In this section, we first evaluate the motion capture framework on the MINIONS dataset. We then conduct ablation studies to evaluate the components of our framework. At last, we evaluate the methods of various tasks, such as 2D-to-3D keypoint lifting and fine-grained video action recognition on our proposed MINIONS dataset.
V-A Implementation Details
We divide the MINIONS dataset into a training set, a validation set, and a testing set by actors. The training set contains 12 actors with 3.2 million frames. The validation set has three actors with 0.9 million frames. The testing set has five actors with 1.4 million frames.
We train and evaluate the SparseNet framework based on different settings on our MINIONS dataset, including IMU-based, monocular vision-based, and multi-modal human motion capture. Following previous works, we apply $\mu_{glo}$ (mean global rotation error), $\sigma_{glo}$ (variance of global rotation error), and MPJPE (Euclidean distance error) to measure the mean error between prediction and ground truth in Euclidean space and angular space. Additionally, we use the Jitter to measure the average jerk of body joints.
During training, we first pre-train the visual branch with a batch size of 64 for 20 epochs. The input frames are resized to $512\times 512$. After that, we train the sparse IMUs branch and post-processing branch with a batch size of 512 for 200 epochs using Adam optimizers with a learning rate of 0.001. All training and test processes run on an NVIDIA GTX A100.
V-B Multi-modal Human Motion Capture
TABLE III: Results on the TotalCapture dataset for motion capture from a monocular camera and full-body IMUs.
Methods TotalCapture [56]
MPJPE $\downarrow$ ($m$) PA-MPJPE $\downarrow$ ($m$)
DiffCap [45] 46.2 29.9
VIP [40] - 26.0
Liu et al. [36] 45.8 -
Ours 36.7 21.6
Comparison. Our experimental results are detailed in Table II. The results from our experiments show that motion capture using the monocular camera is frequently hampered by instability resulting from blurriness or occlusion and leads to large Jitter. In contrast, motion capture systems that rely on IMUs are prone to larger errors in global rotations and Euclidean distance, attributable to the inherent unreliability of the IMUs. However, an integrated system combining just four to six IMUs with a monocular camera can successfully ensure stable motion capture with minimal error and virtually no jitter. Experimental results show that a motion capture system equipped with a monocular camera and more than eight IMUs is unnecessary. They provide negligible improvements in global rotation error and Euclidean distance error, while leading to increased jitter and equipment cost. Furthermore, we conduct experiments for motion capture from a monocular camera and full-body IMUs on the TotalCapture [56] dataset in Table III. Our method outperforms other methods in terms of performance and demonstrates greater flexibility, allowing for adaptation to various sensor configurations.
@@ V-B Multi-modal Human Motion Capture
Visualization. To facilitate a more intuitive comparison, we provide visualization results of vision-based, IMUs-based, and multi-modal motion capture in Figure 8 (c). The vertexes of the 3D human body are colored by the distances between the prediction and the ground truth positions. Additionally, we present the global pose error and positional translation for a motion capture sequence in Figure 8 (a) and (b). The results reveal that the motion capture based on IMUs tends to exhibit a gradual increase in both global pose error and global positional translation over time. However, the integration of a monocular camera and the IMUs can significantly reduce global pose error and also mitigate the inherent global positional drift caused by IMUs.
V-C Benchmarks on other Tasks
@@
In this subject, we evaluate mainstream baselines and benchmarks on tasks including 2D-to-3D pose estimation [42] and Fine-grained Video Action Recognition [54] , which facilitate more comprehensive explorations of our dataset.
TABLE IV: 2D-to-3D pose estimation. #Num: Receptive Field (frames).
Method #Num MPJPE $\downarrow$ ($m$) PA-MPJPE $\downarrow$ ($m$)
@@
2D-to-3D pose estimation. The 2D-to-3D pose estimation task aims to accurately predict the location of 3D joints from the detected 2D human joints, which is a specific type of 3D pose estimation task. The challenge is that depth information is lost in the projection from 3D joints to 2D joints, making the reverse process of deducing 3D joints from 2D observations ill-posed. In TABLE IV, we evaluate the accuracy of the mainstream methods. MPJPE is applied for measuring the average Euclidean distance between the ground truth and predicted joint positions, while PA-MPJPE is a similar metric that first aligns the predicted pose to the ground truth using a rigid transformation before computing the error. To explore the upper bounds of current mainstream methods, we use the groundtruth of 2D joints as input. The experimental results indicate that there is still room for improvement in this task.
@@
Fine-grained Video Action Recognition. Our dataset provides fine-grained action labels and motion videos from various views, which can be validated for applications in fine-grained video action recognition. As shown in TABLE V, we test the accuracy of the mainstream video action understanding methods on our dataset. Following these methods, Top-1 and Top-5 are used to evaluate model performance. Top-1 is the model’s accuracy in predicting the most likely action, while Top-5 is the accuracy at which the true action is within the model’s top-five predictions. Compared to the mainstream video understanding datasets, e.g. Kinetics400 [7], NTU RGB+D 60 [50], and NTU RGB+D 120 [35], our dataset is more challenging since existing methods easily achieve Top-1 accuracy of more than 85% on the Kinetics400 dataset.
TABLE V: Fine-grained Video Action Recognition on datasets.
Method MINIONS (Ours) Kinetics400 [7]
Top-1 $\uparrow$ Top-5 $\uparrow$ Top-1 $\uparrow$ Top-5 $\uparrow$
TSN [59] 44.94% 77.00% 75.89% 92.07%
R(2+1)D [55] 55.48% 85.78% 75.46% 92.28%
SlowFast [18] 53.93% 83.82% 77.03% 92.99%
TSM [34] 59.37% 88.64% 75.12% 91.55%
TPN [63] 64.57% 92.11% 76.74% 92.57%
VideoMAE [58] 73.75% 96.01% 85.30% 96.70%
UniFormerV2 [32] 75.88% 96.87% 89.50% 98.40%
VI Conclusion